Shlomo travels 140 miles in 3 1/2 hours. How fast is he traveling
Answer by jim_thompson5910(35256)

3 1/2 hours = 3.5 hours


d = r*t

140 = r*3.5

140/3.5 = r

40 = r

r = 40


He is traveling 40 miles per hour.